# In-plane transverse polarization in heavy-ion collisions

### v1

@@ Abstract
We give an analytical expression for the last component of the spin polarization $P^{x}$, the in-plane polarization, in heavy-ion collisions that has, to our knowledge, not been discussed in theories nor measured in heavy-ion collision experiments. We also carry out a numerical study of $P^{x}$ using a hydrodynamic model simulation as a cross-check for the analytical formula. It is found that if the temperature-gradient contribution is neglected the simulation result for $P^{x}$ qualitatively agrees with the analytical one. The prediction of $P^{x}$ can be tested in experiments and will contribute to provide a complete and consistent picture of spin phenomena in heavy-ion collisions.
@@ I Introduction
In addition to the global spin polarization, the spin polarization along the beam direction was also proposed in hydrodynamic and transport models (Becattini:2017gcx, ; Xia:2018tes, ) with the expected behavior $P^{z}\sim-\sin(2\phi_{p})$ where $\phi_{p}$ represents the transverse azimuthal angle of the hyperon’s momentum in the reaction plane. However the experimental measurement (STAR:2019erd, ) show an opposite sign behaviour $P^{z}\sim\sin(2\phi_{p})$. The first theoretical explanation of this discrepancy was provided in Ref. (Wu:2019eyi, ; Wu:2020yiz, ), where it was found that the temperature vorticity qualitatively accounts for spin polarization along the beam direction as well as the global spin polarization as a funnction of $\phi_{p}$. Later on, it was found that the contribution from the shear stress tensor can also yield the correct sign for longitudinal polarization (Fu:2021pok, ; Becattini:2021iol, ; Yi:2021ryh, ; Florkowski:2021xvy, ; Wagner:2022gza, ; Wu:2022mkr, ). More recently, it has been proposed that in a thermal model the projected thermal vorticity along with dissipative corrections can describe the behavior of longitudinal polarization (Banerjee:2024xnd, ). Very recently dissipative relativistic spin hydrodynamics was developed from quantum kinetic theory for massive particles with non-local collisions (Weickgenannt:2020aaf, ; Weickgenannt:2021cuo, ; Weickgenannt:2022zxs, ) and gave a good description of longitudinal polarization data (Sapna:2025yss, ). These theoretical approaches collectively indicate that global equilibrium has not been achieved, necessitating the inclusion of off-equilibrium effects in the analysis (Sheng:2021kfc, ).
@@
Inspired by the simple and intuitive blast wave picture of the longitudinal spin polarization, we had performed a comprehensive analysis of spin observables in the framework of the modified or extended blast wave model (Retiere:2003kf, ; Jaiswal:2015saa, ; Yang:2016rnw, ; Yang:2020oig, ; Yang:2022yxa, ). We found analytical expressions or solutions for the longitudinal and global spin polarizations as functions of particle’s momentum and collision centrality under flow-momentum correspondence (Arslan:2024dwi, ). From the analytic solutions, one could clearly see that the global spin polarization is driven by the directed flow, while the longitudinal spin polarization is driven by the ellipticity in flow. The analytical solutions can be improved systematically by perturbative expansion in the small deviations from the flow-momentum correspondence.
In this paper, we provide an analytical expression for the last component of the spin vector $P^{x}$, the in-plane spin polarization, that has, to our knowledge, not been discussed in theories nor measured in experiments. We also carry out a numerical study of $P^{x}$ using a hydrodynamic simulation as a cross-check for such an analytical solution. The prediction of $P^{x}$ can be tested in experiments and may provide a complete and consistent picture for spin phenomena in heavy-ion collisions.
@@ II Model description
$S(x,p)=m_{T}\cosh(\eta-Y)\delta(\tau-\tau_{f})\Theta(R-r)f(p\cdot u).$ (5)
@@
$=$ $\frac{\int_{0}^{R}dr\int_{-\infty}^{\infty}d\eta\int_{0}^{2\pi}d\phi_{s}\;rm_{T}\cosh(\eta-Y)\exp(-\beta p\cdot u)\hat{O}(x,p)}{\int_{0}^{R}dr\int_{-\infty}^{\infty}d\eta\int_{0}^{2\pi}d\phi_{s}\;rm_{T}\cosh(\eta-Y)\exp(-\beta p\cdot u)},$ (6)
where $\hat{O}(x,p)$ is the observable in phase space and $d^{4}x=\tau rd\tau d\eta drd\phi_{s}$. The observable’s spectra in some variables of $p_{T}$, $\phi_{p}$ and $Y$ can be obtained by integration over the rest of the variables in both numerator and denominator in Eq. (6).
@@ III Perturbation method and flow-momentum correspondence
We see that $\delta\rho$ depends on $\Delta\eta$, $Y$, $\Delta\phi$ and $\phi_{p}$ and is of $O(\epsilon)$. The integral measure for $\phi_{s}$ in Eq. (6) can be converted to that for $\phi_{b}$ to $O(\epsilon)$ as
@@
We note that the expansion in Eq. (7) is only valid inside the integral since for large $\bar{\rho}$ and $\Delta\eta$, the $\delta\rho$ terms can easily become larger than 1. When integrated over $r$ and $\Delta\eta$, the exponential factor suppresses the contribution from large $\Delta\eta$ ($\cosh\bar{\rho}$ and $\sinh\bar{\rho}$ are finite since $\bar{\rho}<1$ for $\rho_{0}\sim 1$).
@@ IV Polarization vectors: general results
$\hat{P}_{\xi}^{\mu}=$ $-\frac{1}{2m}\epsilon^{\mu\nu\sigma\tau}(1-f)\frac{p_{\tau}p^{\rho}}{E_{p}}\hat{t}_{\nu}\xi_{\rho\sigma},$ (15)
where the vector $\hat{t}$ is chosen to be $\hat{t}^{\mu}=(1,0,0,0)$ corresponding to the normal direction of the freeze-out hyper-surface for $\eta=Y=0$, $\omega^{\mu\nu}$ and $\xi^{\mu\nu}$ denote the thermal vorticity and thermal shear stress tensors respectively defined as
@@
$\times\sum_{n=0}\left(m_{T}C_{z,n}^{\omega}+p_{T}C_{z,n}^{\xi}\right)K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}),$ (18)
@@
$\times\sum_{n=0}\left(m_{T}C_{z,n}^{\omega}+p_{T}C_{z,n}^{\xi}\right)K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}).$ (20)
@@
$\times K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho}),$ (21)
@@ V Transverse polarizations
We see from Eqs. (18,20,21) that both $P^{x}$ and $P^{y}$ are driven by the directed flow. The $\phi_{p}$ pattern of $P^{x}$ is $\sin(2\phi_{p})$, same as $P^{z}$ but with a slightly smaller magnitude which will be shown in Fig. 7. There are two terms in $P^{y}(\phi_{p})$, the first term would be vanishing if we take $\Delta\phi=0$ in $\hat{P}^{y}(\phi_{p})$ before the integration over $\Delta\phi$ in calculating its expectation value as in Ref. (Arslan:2024dwi, ), while the second term is proportional to $P^{x}(\phi_{p})$ up to a modulation factor $\sim\sin\phi_{p}$. We can combine $P^{x}(\phi_{p})$ and $P^{y}(\phi_{p})$ to form a vector in the transverse plane
@@ VI.1 Numerical results from analytical formulas
Figure 4: The results for $P^{z}$ (right panel) and $P_{H}\equiv-P^{y}$ (left panel) as functions of $\phi_{p}$ following Eq.(20). The black solid lines represent the calculated results based on analytical formulas.
@@
The calculated results for the transverse momentum dependence of polarization following the analytical formulas in Eq. (21). The results are shown in Fig. 5. The data show that $P_{\sin 2\phi}^{z}(p_{T})$ is almost a constant when $p_{T}\gtrsim 1\text{ GeV}$ and this behavior can be well described by the analytical formula in Eq. (21). In contrast, the result based on the analytical formula in Ref. (Arslan:2024dwi, ) grows with increasing $p_{T}$. The improved results obtained in this work therefore indicate that relativistic formulation is important in describing $P_{\sin 2\phi}^{z}(p_{T})$ correctly.
@@ VI.2 Simulation results by hydrodynamical models
The first two terms are the velocity-gradient terms in Eq. (14) and Eq. (15). The last term, $\hat{P}_{T}^{\mu}$, represents the polarization caused by the temperature gradient ($T$-gradient) and is given by:
@@ Acknowledgements.
A large part of this work was completed during the time when Q.W. was a visiting professor in the nuclear theory group at McGill University. The authors thank X.-L. Sheng and S. Lin for helpful discussion. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 12135011, and in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [SAPIN-2020-00048, SAPIN-2024-00026].

### v2

@@ Abstract
We give an analytical expression for the in-plane polarization $P^{x}$, in heavy-ion collisions that has, to our knowledge, not been measured in heavy-ion collision experiments. We also carry out a numerical study of $P^{x}$ using a hydrodynamic model simulation as a cross-check for the analytical formula. It is found that if the temperature-gradient contribution is neglected the simulation result for $P^{x}$ qualitatively agrees with the analytical one. The prediction of $P^{x}$ can be tested in experiments and will contribute to provide a complete and consistent picture of spin phenomena in heavy-ion collisions.
@@ I Introduction
In addition to the global spin polarization, the spin polarization along the beam direction was also proposed in hydrodynamic and transport models (Becattini:2017gcx, ; Xia:2018tes, ) with the expected behavior $P^{z}\sim-\sin(2\phi_{p})$ where $\phi_{p}$ represents the transverse azimuthal angle of the hyperon’s momentum in the reaction plane. However the experimental measurement (STAR:2019erd, ) shows an opposite sign behaviour $P^{z}\sim\sin(2\phi_{p})$. The first theoretical explanation of this discrepancy was provided in Ref. (Wu:2019eyi, ; Wu:2020yiz, ), where it was found that the temperature vorticity qualitatively accounts for spin polarization along the beam direction as well as the global spin polarization as a function of $\phi_{p}$. Later on, it was found that the contribution from the shear stress tensor can also yield the correct sign for longitudinal polarization (Fu:2021pok, ; Becattini:2021iol, ; Yi:2021ryh, ; Florkowski:2021xvy, ; Wagner:2022gza, ; Wu:2022mkr, ). More recently, it has been proposed that in a thermal model the projected thermal vorticity along with dissipative corrections can describe the behavior of longitudinal polarization (Banerjee:2024xnd, ). Very recently dissipative relativistic spin hydrodynamics was developed from quantum kinetic theory for massive particles with non-local collisions (Weickgenannt:2020aaf, ; Weickgenannt:2021cuo, ; Weickgenannt:2022zxs, ) and gave a good description of longitudinal polarization data (Sapna:2025yss, ). These theoretical approaches collectively indicate that global equilibrium has not been achieved, necessitating the inclusion of off-equilibrium effects in the analysis (Sheng:2021kfc, ).
@@
Inspired by the simple and intuitive blast wave picture of the longitudinal spin polarization, we had performed a comprehensive analysis of spin observables in the framework of the modified or extended blast wave model (Retiere:2003kf, ; Jaiswal:2015saa, ; Yang:2016rnw, ; Yang:2020oig, ; Yang:2022yxa, ). We found analytical expressions or solutions for the longitudinal and global spin polarizations as functions of particle’s momentum and collision centrality under flow-momentum correspondence (Arslan:2024dwi, ). From the analytic solutions, one could clearly see that the global spin polarization is driven by the directed flow, while the longitudinal spin polarization is driven by the ellipticity in the emission zone and flow. The analytical solutions can be improved systematically by perturbative expansion in the small deviations from the flow-momentum correspondence.
In this paper, we provide an analytical expression for the in-plane spin polarization $P^{x}$, that has, to our knowledge, not been measured in experiments. There are some earlier theoretical results for $P^{x}$ from hydrodynamical or transport model simulations (Karpenko:2016jyx, ; Xia:2018tes, ; Sun:2021nsg, ), but the magnitudes of $P^{x}$ are very small compared with those of $P^{y}$. Our analytical formulas show that $P^{x}$ and $P^{y}$ have the same order of magnitudes. We also carry out a numerical study of $P^{x}$ using a hydrodynamic simulation as a cross-check for such an analytical solution. The prediction of $P^{x}$ can be tested in experiments and may provide a complete and consistent picture for spin phenomena in heavy-ion collisions.
@@ II Model description
$S(x,p)=$ $m_{T}\cosh(\eta-Y)\delta(\tau-\tau_{f})\Theta(1-\tilde{r})f(p\cdot u)$
$\approx$ $m_{T}\cosh(\eta-Y)\delta(\tau-\tau_{f})\Theta(R-r)f(p\cdot u)$
$-\frac{\epsilon}{2}\cos(2\phi_{s})m_{T}\cosh(\eta-Y)\delta(\tau-\tau_{f})\delta(1-r/R)f(p\cdot u),$ (5)
where the last term in the second equality originates from the elliptic deformation of the emission zone characterized by the normalized elliptic radius Retiere:2001ed defined as
$\tilde{r}=\sqrt{\left(\frac{r}{R_{x}}\right)^{2}\cos^{2}\phi_{s}+\left(\frac{r}{R_{x}}\right)^{2}\sin^{2}\phi_{s}}\;.$ (6)
@@
$=$ $\frac{\int_{0}^{R}dr\int_{-\infty}^{\infty}d\eta\int_{0}^{2\pi}d\phi_{s}\;rm_{T}\cosh(\eta-Y)\exp(-\beta p\cdot u)\hat{O}(x,p)}{\int_{0}^{R}dr\int_{-\infty}^{\infty}d\eta\int_{0}^{2\pi}d\phi_{s}\;rm_{T}\cosh(\eta-Y)\exp(-\beta p\cdot u)}$
$-\frac{\epsilon R^{2}\int_{-\infty}^{+\infty}d\eta\int_{0}^{2\pi}\cos(2\phi_{s})d\phi_{s}m_{T}\cosh(\eta-Y)\exp(-\beta p\cdot u)\hat{O}(x,p)|_{r=R}}{2\int_{0}^{R}dr\int_{-\infty}^{\infty}d\eta\int_{0}^{2\pi}d\phi_{s}\;rm_{T}\cosh(\eta-Y)\exp(-\beta p\cdot u)},$ (7)
where $\hat{O}(x,p)$ is the observable in phase space and $d^{4}x=\tau rd\tau d\eta drd\phi_{s}$. The observable’s spectra in some variables of $p_{T}$, $\phi_{p}$ and $Y$ can be obtained by integration over the rest of the variables in both numerator and denominator in Eq. (7). Note that the observables we consider in this paper are all up to $O(\epsilon)$, so the $O(\epsilon)$ contribution in the denominators in Eq. (7) can be neglected since the numerators are already up to $O(\epsilon)$.
@@ III Perturbation method and flow-momentum correspondence
We see that $\delta\rho$ depends on $\Delta\eta$, $Y$, $\Delta\phi$ and $\phi_{p}$ and is of $O(\epsilon)$. Applying the identity $\tan(\phi_{b})=(R_{x}^{2}/R_{y}^{2})\tan(\phi_{s})$, the integral measure for $\phi_{s}$ in Eq. (7) can be converted to that for $\phi_{b}$ to $O(\epsilon)$ as
@@
We note that the expansion in Eq. (8) is only valid inside the integral since for large $\bar{\rho}$ and $\Delta\eta$, $\mathcal{O}(\epsilon)$ terms can easily become larger than 1. When integrated over $r$ and $\Delta\eta$, the exponential factor suppresses the contribution from large $\Delta\eta$ ($\cosh\bar{\rho}$ and $\sinh\bar{\rho}$ are finite since $\bar{\rho}<1$ for $\rho_{0}\sim 1$).
@@ IV Polarization vectors: general results
$\hat{P}_{\xi}^{\mu}=$ $-\frac{1}{2m}\epsilon^{\mu\nu\sigma\tau}(1-f)\frac{p_{\tau}p^{\rho}}{p\cdot\hat{t}}\hat{t}_{\nu}\xi_{\rho\sigma},$ (16)
where the vector $\hat{t}$ is chosen to be $\hat{t}^{\mu}=(1,0,0,0)$ corresponding to the normal direction of the freeze-out hyper-surface $T_{f}=T(\tau_{f})$ for $\eta=Y=0$, $\omega^{\mu\nu}$ and $\xi^{\mu\nu}$ denote the thermal vorticity and thermal shear stress tensors defined respectively as
@@
Here we make a few remarks about the freeze-out hyper-surface and the $T$-gradient contribution. In the blast wave model it is normally assumed that $T$ is only a function of the proper time $\tau$, i.e. $T=T(\tau)$, so the $T$-gradient contribution only comes from $\partial T/\partial\tau$. However one can show that there are non-vanishing $\partial T/\partial\tau$ terms in $P_{\omega}$ and $P_{\xi}$, but these terms cancel out in $P_{\omega}+P_{\xi}$ when $\eta=Y=0$ and $\hat{t}^{\mu}=(1,0,0,0)$ (the normal direction of the freeze-out hyper-surface at $\eta=Y=0$). On the other hand, it can be shown in quantum statistical theory that if $\hat{t}^{\mu}$ is chosen to be the normal direction of the isothermal freeze-out hyper-surface, the $T$-gradient contribution is vanishing to any order of gradient expansion (Sheng:2025cjk, ).
@@
$\times\sum_{n=0}\left(m_{T}C_{z,n}^{\omega}+p_{T}C_{z,n}^{\xi}\right)K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho})$
$+\frac{\epsilon R}{16mT}\sin(2\phi_{p})\frac{1}{N_{0}}\int_{0}^{p_{T}^{\mathrm{max}}}dp_{T}K_{1}\left(\beta m_{T}\cosh\rho_{0}\right)$
$\times\left\{p_{T}^{3}\left(\sinh\rho_{0}-\rho_{0}\cosh\rho_{0}\right)\right.\left[I_{0}\left(\beta p_{T}\sinh\rho_{0}\right)-I_{4}\left(\beta p_{T}\sinh\rho_{0}\right)\right]$
$\left.+2p_{T}^{2}m_{T}\rho_{0}\sinh\rho_{0}\left[I_{1}\left(\beta p_{T}\sinh\rho_{0}\right)-I_{3}\left(\beta p_{T}\sinh\rho_{0}\right)\right]\right\},$ (19)
@@
$\times\sum_{n=0}\left(m_{T}C_{z,n}^{\omega}+p_{T}C_{z,n}^{\xi}\right)K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho})$
$+\frac{\epsilon R}{32mT}\frac{1}{N_{0}}\int_{0}^{p_{T}^{\mathrm{max}}}dp_{T}K_{1}\left(\beta m_{T}\cosh\rho_{0}\right)$
$\times\left\{p_{T}^{3}\left(\sinh\rho_{0}-\rho_{0}\cosh\rho_{0}\right)\right.\left[I_{0}\left(\beta p_{T}\sinh\rho_{0}\right)-I_{4}\left(\beta p_{T}\sinh\rho_{0}\right)\right]$
$\left.+2p_{T}^{2}m_{T}\rho_{0}\sinh\rho_{0}\left[I_{1}\left(\beta p_{T}\sinh\rho_{0}\right)-I_{3}\left(\beta p_{T}\sinh\rho_{0}\right)\right]\right\}.$ (21)
@@
$\times K_{1}(\beta m_{T}\cosh\bar{\rho})I_{n}(\beta p_{T}\sinh\bar{\rho})$
$+\frac{\epsilon R}{32mT}\frac{1}{N_{0}(p_{T})}K_{1}\left(\beta m_{T}\cosh\rho_{0}\right)$
$\times\left\{p_{T}^{2}\left(\sinh\rho_{0}-\rho_{0}\cosh\rho_{0}\right)\right.\left[I_{0}\left(\beta p_{T}\sinh\rho_{0}\right)-I_{4}\left(\beta p_{T}\sinh\rho_{0}\right)\right]$
$\left.+2p_{T}m_{T}\rho_{0}\sinh\rho_{0}\left[I_{1}\left(\beta p_{T}\sinh\rho_{0}\right)-I_{3}\left(\beta p_{T}\sinh\rho_{0}\right)\right]\right\},$ (22)
@@ V Transverse polarizations
We see from Eqs. (19,21,22) that both $P^{x}$ and $P^{y}$ are driven by the directed flow. The $\phi_{p}$ pattern of $P^{x}$ is $\sin(2\phi_{p})$, same as $P^{z}$ but with a slightly smaller magnitude which will be shown in Fig. 7. There are two terms in $P^{y}(\phi_{p})$, the first term would be vanishing if we take $\Delta\phi=0$ in $\hat{P}^{y}(\phi_{p})$ before the integration over $\Delta\phi$ in calculating its expectation value as in Ref. (Arslan:2024dwi, ), while the second term is proportional to $P^{x}(\phi_{p})$ up to a modulation factor $\sim\sin\phi_{p}$. We can combine $P^{x}(\phi_{p})$ and $P^{y}(\phi_{p})$ to form a vector in the transverse plane
@@ VI.1 Numerical results from analytical formulas
Figure 4: The results for $P^{z}$ (left panel) and $P_{H}\equiv-P^{y}$ (right panel) as functions of $\phi_{p}$ following Eq.(21). The black solid lines represent the calculated results based on analytical formulas.
@@
The calculated results for the transverse momentum dependence of polarization following the analytical formulas in Eq. (22). In contrast to the analytical formula in Arslan:2024dwi , Eq. (22) fits better to the data.
@@ VI.2 Simulation results by hydrodynamical models
The first two terms are the velocity-gradient terms in Eq. (15) and Eq. (16) with $\hat{t}^{\mu}=u^{\mu}$. The last term, $\hat{P}_{T}^{\mu}$, represents the polarization caused by the temperature gradient ($T$-gradient) and is given by:
@@ Acknowledgements.
A large part of this work was completed during the time when Q.W. was a visiting professor in the nuclear theory group at McGill University. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 12135011, and in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) [SAPIN-2020-00048, SAPIN-2024-00026].
Appendix A The coefficients in Eq. (19)
The explicit forms of $C^{\omega}_{z,n}$ with $n=0,1,\cdots,4$ in Eq. (19) are given as
$C_{z,0}^{\omega}=$ $-\frac{1}{2R^{2}}\rho_{0}\beta\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}^{2}\sinh\bar{\rho}\cosh\bar{\rho},$
$C_{z,1}^{\omega}=$ $-\frac{1}{R^{2}}\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}\cosh\bar{\rho}-\frac{1}{R}\left(3\rho_{2}+\frac{5}{2}\epsilon\rho_{0}\right)p_{T}\sinh\bar{\rho}$
$+\frac{1}{R^{2}}\beta\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}m_{T}\sinh^{2}\bar{\rho},$
$C_{z,2}^{\omega}=$ $4\rho_{2}\frac{1}{R}m_{T}\cosh\bar{\rho}-2\epsilon\frac{1}{r}m_{T}\sinh\bar{\rho},$
$C_{z,3}^{\omega}=$ $\frac{1}{R^{2}}\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}\cosh\bar{\rho}+\frac{1}{R}\left(-\rho_{2}+\frac{5}{2}\epsilon\rho_{0}\right)p_{T}\sinh\bar{\rho}$
$-\frac{1}{R^{2}}\beta\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}m_{T}\sinh^{2}\bar{\rho},$
$C_{z,4}^{\omega}=$ $\frac{1}{2R^{2}}\rho_{0}\beta\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}^{2}\sinh\bar{\rho}\cosh\bar{\rho}.$ (30)
The explicit forms of $C^{\xi}_{z,n}$ with $n=0,1,\cdots,5$ are given as
$C_{z,0}^{\xi}=$ $\frac{1}{R}2\left(\rho_{2}+\epsilon\rho_{0}\right)p_{T}\cosh\bar{\rho}+\frac{1}{R^{2}}\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}\sinh\bar{\rho}$
$+\epsilon\frac{1}{r}p_{T}\sinh\bar{\rho}-\frac{3r}{2R^{2}}\beta\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)p_{T}m_{T}\cosh\bar{\rho}\sinh\bar{\rho}$
$+\frac{1}{R}\beta\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)m_{T}p_{T}\sinh^{2}\bar{\rho},$
$C_{z,1}^{\xi}=$ $-\frac{1}{R^{2}}\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rm_{T}\cosh\bar{\rho}-\frac{1}{R}\left(3\rho_{2}+\frac{5}{2}\epsilon\rho_{0}\right)m_{T}\sinh\bar{\rho}$
$+\frac{1}{R^{2}}\beta\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}^{2}\cosh^{2}\bar{\rho}$
$+\frac{1}{R^{2}}\beta\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rm_{T}^{2}\sinh^{2}\bar{\rho}$
$-\frac{1}{R}\beta\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)p_{T}^{2}\cosh\bar{\rho}\sinh\bar{\rho},$
$C_{z,2}^{\xi}=$ $0,$
$C_{z,3}^{\xi}=$ $\frac{1}{R^{2}}\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rm_{T}\cosh\bar{\rho}+\frac{1}{R}\left(-\rho_{2}+\frac{5}{2}\epsilon\rho_{0}\right)m_{T}\sinh\bar{\rho}$
$-\frac{1}{2R^{2}}\beta\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}^{2}\cosh^{2}\bar{\rho}$
$-\frac{1}{R^{2}}\beta\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rm_{T}^{2}\sinh^{2}\bar{\rho}$
$+\frac{1}{2R}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)\beta p_{T}^{2}\cosh\bar{\rho}\sinh\bar{\rho},$
$C_{z,4}^{\xi}=$ $\frac{1}{R}2\left(\rho_{2}-\epsilon\rho_{0}\right)p_{T}\cosh\bar{\rho}-\frac{1}{R^{2}}\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}\sinh\bar{\rho}$
$+\epsilon\frac{1}{r}p_{T}\sinh\bar{\rho}-\frac{1}{R}\beta\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)m_{T}p_{T}\sinh^{2}\bar{\rho}$
$+\frac{3}{2R^{2}}\beta\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}m_{T}\cosh\bar{\rho}\sinh\bar{\rho},$
$C_{z,5}^{\xi}=$ $-\frac{1}{2R^{2}}\beta\rho_{0}\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)rp_{T}^{2}\cosh^{2}\bar{\rho}$
$+\frac{1}{2R}\beta\left(\rho_{2}+\frac{1}{2}\epsilon\rho_{0}\right)p_{T}^{2}\cosh\bar{\rho}\sinh\bar{\rho}.$ (31)